OpenStudy (anonymous):

IF ABC IS an acute triangle show that:

OpenStudy (anonymous):

\[\sin \frac{ A }{ 2 }\sin \frac{ B }{ 2 }\sin \frac{ C }{ 2 }\le \frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

@eliassaab @thomaster @undeadknight26 @ganeshie8 @

OpenStudy (anonymous):

Here is a proof with Mathematica

OpenStudy (anonymous):

The maximum 1/8 is attained when A=B=C=60 degrees

OpenStudy (anonymous):

The way to do it properly is the following. Notice that \[ f(x) = \ln(\sin(x)) \] is concave on the interval \([0, \pi] \)

OpenStudy (anonymous):

Hence \[ \frac 1 3 f(A/2) +\frac 1 3 f(B/2)+ \frac 1 3 f(C/2)\le f\left( \frac 1 3 \frac A 2 +\frac 1 3 \frac B 2 +\frac 1 3 \frac C 2 \right)=\\ f\left( \frac 1 3 \left( \frac A 2 +\frac B 2 +\frac C 2 \right) \right)=\\ f\left( \frac 1 3 \left( \frac \pi 2 \right) \right)=f\left( \frac \pi 6 \right)= \ln(\sin(\pi/6)=\ln(1/2)\\ \] The rest is easy

OpenStudy (anonymous):

\[ \frac 1 3 f(A/2) +\frac 1 3 f(B/2)+ \frac 1 3 f(C/2)=\\ \frac 1 3 \ln( \sin((A/2)) +\frac 1 3 \ln(\sin(B/2))+ \frac 1 3 \ln\sin((C/2))=\\ \frac 1 3 \left(\ln( \sin((A/2)) + \ln(\sin(B/2))+ \ln\sin((C/2))\right)=\\ \frac 1 3 \left(\ln( \sin((A/2) \sin((B/2)\sin((C/2)) \right)=\\ \ln( \sin((A/2) \sin((B/2)\sin((C/2))^{1/3}\le \ln(1/2)\\ \sin((A/2) \sin((B/2)\sin((C/2))^{1/3}\le \frac 12 \\ \sin((A/2) \sin((B/2)\sin((C/2))\le \left(\frac 12\right)^3=\frac 1 8 \\ \]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Actually this exercise is strictly related to jessner theorem

ganeshie8 (ganeshie8):

beautiful ! another way using lagrange multipliers : \(\large f : A+ B + C = \pi \) \(\large g : \sin \frac{A}{2} \sin \frac{B}{2}\sin \frac{C}{2}\) \(\large \nabla g = \lambda \nabla f\) solving gives the max value of \(\large g= \frac{1}{8}\) at \(\large A = B = C = \frac{\pi}{3}\)